y-x=7
y+x=9
is it one, no, or infinitely solutions?

Answer by jim_thompson5910(35256)

Add the two equations

y-x=7
y+x=9
------
2y+0x = 16


So 2y = 16 which means y = 8. I divided both sides by 2 to isolate y.

Because y = 8, we can use this to find x

y-x=7

8-x=7

8 = 7+x

8 - 7 = x

1 = x

x = 1

So we get one unique solution and it is (x,y) = (1,8)